1. A jet airplane in level flight has a mass of 8.65 ✕ 104 kg, and the two wings have an estimated total area of 92.0 m2.

(a) What is the pressure difference between the lower and upper surfaces of the wings?

(b) If the speed of air under the wings is 226 m/s, what is the speed of the air over the wings? Assume air has a density of 1.29 kg/m3.

2. A long steel I-beam used in bridge construction has a length of 20.0 m. The beam may be exposed to temperatures as low as −21°C in the winter and as high as 38°C

in the summer. What is the difference in length of the beam between these two temperature extremes? (Give the absolute value of the difference. Enter your answer in mm.)
Answer #1

1) Area , A = 92 m^2

m = 8.65 *10^4 Kg

let the pressure difference is P

P * Area = m * g

P * 92 = 8.65 *10^4 * 9.8

P = 9214 Pa

the pressure difference is 9214 Pa

b) let the speed at the top is v

for the pressure difference is P

P = 0.50 * density * (v^2 - 226^2)

9214 = 0.50 * 1.29 * (v^2 - 226^2)

v = 255.6 m/s

the speed of the air over the wings is 255.6 m/s